# DeFi Arbitrage in Hedged Liquidity Tokens

### v2

@@ 4.1 Implied Volatility
First, consider the case with zero risk-free rate $r=0$. We note that $\lim_{\sigma\searrow 0}G(\sigma)=0$ (with $\lim_{\sigma\searrow 0}G^{\prime}(\sigma)=-\hat{\gamma}\sqrt{\frac{\Delta t}{2\pi}}$) and $\lim_{\sigma\nearrow\infty}G(\sigma)=2$. Therefore, there exists at least one implied volatility $\sigma^{I}>0$. Additionally, $\bar{\sigma}:=\frac{\hat{\gamma}}{2+\hat{\gamma}}\sqrt{\frac{8}{\pi\Delta t}}$ is the unique positive root of $G^{\prime}$. Therefore, there must exist a unique implied volatility $\sigma^{I}>\bar{\sigma}$.
Assume, now, a strictly positive risk-free rate $r>0$. First, we note that $\lim_{\sigma\searrow 0}G(\sigma)=2(1-e^{-\frac{1}{2}r\Delta t})-\hat{\gamma}(e^{-\frac{1}{2}r\Delta t}-e^{-r\Delta t})>0$ and $\lim_{\sigma\nearrow\infty}G(\sigma)=2$. Therefore, there exists an implied volatility $\sigma^{I}>0$ if, and only if, $\inf_{\sigma>0}G(\sigma)\leq 0$. There exists a unique root $G^{\prime}(\bar{\sigma})=0$ if, and only if, $\Delta t\leq\overline{\Delta}t$ (otherwise $G^{\prime}(\sigma)>0$ for every $\sigma>0$) with $\bar{\sigma}$ given in the statement of this lemma. Assume $\Delta t\leq\overline{\Delta}t$, then since this root $\bar{\sigma}$ is unique, it must follow that $\inf_{\sigma>0}G(\sigma)\in\{G(\bar{\sigma}),\lim_{\sigma\searrow 0}G(\sigma)\}$. From this, and recalling the relation between the sign of $G$ and the risk-neutral valuation, the result trivially follows. ∎
@@
Before continuing, we want to provide a numerical example to demonstrate all three possible outcomes for the set of implied volatilities. We will do this by considering the Polygon blockchain with different fee levels $\gamma$ to demonstrate the different possible settings for the implied volatility. In each of these cases, $\Delta t<\overline{\Delta}t$ by orders of magnitude.
@@
Consider a constant product market maker on the Polygon blockchain ($\Delta t=2$ seconds). Take the risk-free rate to be $r=5\%$ (annualized). For these examples, recall that the risk-neutral valuation and other formulas employed throughout this work utilize the realized fee level $\hat{\gamma}=\gamma/(1-\gamma)$. For any choice of $\gamma$, there are two possible outcomes. The less interesting possibility is that $V_{0}(P)<2\sqrt{P}$ so that depositing liquidity for the price of $2\sqrt{P}$ is expected to lose value, i.e., the expected discounted value of the fees would not cover this initial deposit value. [^5] Within this work, we have assumed that it is not possible to sell liquidity tokens short and, therefore, no arbitrage opportunity exists if $V_{0}(P)<2\sqrt{P}$. The more interesting scenario is where $V_{0}(P)\geq 2\sqrt{P}$ which (with a strict inequality) results in an arbitrage opportunity since the purchase price of the liquidity tokens $2\sqrt{P}$ is below the risk-neutral valuation of the fees. Therefore, in this latter scenario, a rational investor can deposit the liquidity for $2\sqrt{P}$ and, with appropriate hedging, obtain the value $V_{0}(P)$. [^6] We refer the interested reader to [15, Chapter 1.4] for an initial discussion on arbitrage.
@@
Figure 4: Example 4.4: The shaded region indicates fee-volatility $(\gamma,\sigma)$ pairs that provide arbitrage opportunities. The marked points indicate the implied volatilities $\sigma^{I}$ at $\gamma\in\{1bps,1.14114bps,5bps\}$ as taken within the example.
  • If $\gamma=1bps$ then $\overline{\Delta}t\approx 8.48$ hours and $\bar{\sigma}\approx 0.3168$ with $\hat{\gamma}^{*}(\bar{\sigma})\approx 1.4962bps>\hat{\gamma}$. Therefore, there does not exist an implied volatility at this fee level, i.e., a risk-neutral investor would never deposit liquidity at the AMM. This is shown in Figure 4 by the dotted line.
  • If $\gamma\approx 1.4114bps$ (so that $\hat{\gamma}\approx 1.4116bps$) then $\overline{\Delta}t\approx 11.97$ hours and $\bar{\sigma}\approx 0.4472$ with $\hat{\gamma}^{*}(\bar{\sigma})=\hat{\gamma}$. Therefore, the unique implied volatility is given by $\sigma^{I}=\bar{\sigma}\approx 0.4472$. Notably, if $\sigma\neq\sigma^{I}$ then a risk-neutral investor would not pool liquidity at the AMM.
@@
  • If $\gamma=5bps$ then $\overline{\Delta}t\approx 42.40$ hours and $\bar{\sigma}\approx 1.5846$ with $\hat{\gamma}^{*}(\bar{\sigma})\approx 2.7002bps<\hat{\gamma}$. Therefore, there exists two implied volatilities: $\sigma^{I}_{1}\approx 0.0644<\bar{\sigma}<\sigma^{I}_{2}\approx 3.1047$. Notably, as proven in Corollary 4.3 and shown Figure 4, if $\sigma\in(\sigma^{I}_{1},\sigma^{I}_{2})$ then $V(P_{0})>2\sqrt{P_{0}}$ and a risk-neutral investor would deposit liquidity at the AMM. In contrast, if $\sigma\in(0,\sigma^{I}_{1})\cup(\sigma^{I}_{2},\infty)$ then $V(P_{0})<2\sqrt{P_{0}}$ and a risk-neutral investor would not pool liquidity at the AMM.
@@
As evidenced in Example 4.4 above (and comparing to the parameters of, e.g., Example 2.3), non-uniqueness of the implied volatility can easily occur in practice. It becomes important to understand which of $\{\sigma^{I}_{1},\sigma^{I}_{2}\}$ should be quoted. Herein, as $\sigma^{I}_{2}$ converges to the unique solution $\sigma^{I}$ when the risk-free rate approaches 0, i.e., $\lim_{r\searrow 0}\sigma^{I}_{2}(r)=\sigma^{I}(0)$ with dependence on the risk-free rate made explicit, we take this upper implied volatility to be the more meaningful setting. In comparison, the lower implied volatility $\sigma^{I}_{1}$ converges to 0 as the risk-free rate approaches 0, i.e., $\lim_{r\searrow 0}\sigma^{I}_{1}(r)=0$.
@@ 4.2 Estimating the Arbitrage-Free Price of a Liquidity Token
In contrast to the market price $2\sqrt{P}$ utilized for the implied volatility in Section 4.1 above, herein we want to calibrate the pricing of the liquidity token to the observed data. To do this, we will first find the volatility $\sigma^{M}>0$ so that we observe martingale pricing in the data. Recall from Theorem 3.2 that for any volatility $\sigma>0$, and with abuse of notation so that $\hat{\gamma}^{*}(\sigma)$ explicitly depends on the volatility, we have
$\frac{2\hat{\gamma}\sqrt{P_{0}}}{\hat{\gamma}^{*}(\sigma)}=e^{-r\Delta t}\mathbb{E}\left[\frac{2\hat{\gamma}\sqrt{P_{\Delta t}}}{\hat{\gamma}^{*}(\sigma)}+\hat{\gamma}F(P_{0},P_{\Delta t})\right].$
Further, following Assumption 3.1, this is equivalent to
$2\hat{\gamma}=2\hat{\gamma}\exp\left(-\frac{1}{2}\left(r+\frac{\sigma^{2}}{4}\right)\Delta t\right)+e^{-r\Delta t}\hat{\gamma}^{*}(\sigma)\mathbb{E}\left[\frac{\hat{\gamma}F(P_{0},P_{\Delta t})}{\sqrt{P_{0}}}\right].$
Herein we now wish to calibrate the volatility $\sigma^{M}$ to the observed data in the CPMM pool, i.e., with the observed collected fees $\bar{f}_{t}:=\hat{\gamma}F(P_{t-\Delta t},P_{t})$ along the observed price process $(P_{t})$. Using $N$ observations, [^7] We note that a large number of blocks may not have any realized trades in any given CPMM. As such no price changes are recorded for those blocks. While this apparently violates the geometric Brownian Motion assumption in Assumption 3.1, we wish to note that this is very similar situation to the one occurring in traditional financial markets where trading is not time homogeneous and occurs mostly at the opening and closing times, see e.g., [20, 12]. we solve the following for $\sigma^{M}$
$2\hat{\gamma}=2\hat{\gamma}\exp\left(-\frac{1}{2}\left(r+\frac{(\sigma^{M})^{2}}{4}\right)\Delta t\right)+\frac{e^{-r\Delta t}\hat{\gamma}^{*}(\sigma^{M})}{N}\sum_{n=1}^{N}\frac{\bar{f}_{n\Delta t}}{\sqrt{P_{(n-1)\Delta t}}}.$ (4.1)
For notational simplicity, throughout this section we take $C:=\frac{e^{-r\Delta t}}{N\hat{\gamma}}\sum_{n=1}^{N}\frac{\bar{f}_{n\Delta t}}{\sqrt{P_{(n-1)\Delta t}}}>0$. Within the following proposition, we provide conditions for the existence and uniqueness of a calibrated volatility $\sigma^{M}$ satisfying (4.1) which provides arbitrage profits from being long in the liquidity tokens at the market price $2\sqrt{P}$. Following these mathematical results, we provide the specific details on the calibration and repricing procedure for the data from Example 2.3.
Corollary 4.5.
Recall the notation of the implied volatility from Lemma 4.2. Let $G_{C}:\mathbb{R}_{++}\to\mathbb{R}$ be defined by:
$G_{C}(\sigma):=C+\left[1-\operatorname{e}^{-\frac{1}{2}\left(r+\frac{\sigma^{2}}{4}\right)\Delta t}\right]-\left[\Phi\left(\frac{\left(r+\frac{\sigma^{2}}{2}\right)\sqrt{\Delta t}}{\sigma}\right)-e^{-r\Delta t}\Phi\left(\frac{\left(r-\frac{\sigma^{2}}{2}\right)\sqrt{\Delta t}}{\sigma}\right)\right].$
The following conditions partition all possible cases for the calibrated volatility $\sigma^{M}>0$, i.e., such that $G_{C}(\sigma^{M})=0$ with $\hat{\gamma}\geq\hat{\gamma}^{*}(\sigma^{M})$.
  • If $r=0$ then:
    • If $\gamma\leq\bar{\gamma}_{*}^{0}:=\frac{2(1-\exp(-1/\pi))}{2\Phi(\sqrt{2/\pi})-\exp(-1/\pi)}\approx 64.32\%$ then:
        there does not exist a calibrated volatility if $G_{C}(\sigma^{I})>0$; and
        there exists a unique calibrated volatility $\sigma^{M}\in(0,\sigma^{I}]$ if $G_{C}(\sigma^{I})\leq 0$.
    • If $\gamma>\bar{\gamma}_{*}^{0}$ then define $\bar{\sigma}_{*}^{0}:=\sqrt{\frac{8}{\pi\Delta t}}$ and:
        there does not exist a calibrated volatility if $G_{C}(\bar{\sigma}_{*}^{0})>0$;
        there exists a unique calibrated volatility $\sigma^{M}=\bar{\sigma}_{*}^{0}$ if $G_{C}(\bar{\sigma}_{*}^{0})=0$;
        there exists a unique calibrated volatility $\sigma^{M}\in(0,\bar{\sigma}_{*}^{0})$ if $G_{C}(\bar{\sigma}_{*}^{0})<0$ and $G_{C}(\sigma^{I})<0$; and
        there exists exactly two distinct calibrated volatilities $\sigma^{M}_{1}<\bar{\sigma}_{*}^{0}<\sigma^{M}_{2}\leq\sigma^{I}$ if $G_{C}(\bar{\sigma}_{*}^{0})<0$ and $G_{C}(\sigma^{I})\geq 0$.
  • If $r>0$ then:
    • If $\Delta t>\overline{\Delta t}$ then no calibrated volatility exists.
    • If $\Delta t\leq\overline{\Delta t}$ then define
      $\bar{\sigma}_{*}$ $:=r\sqrt{\frac{\Delta t}{-W\left(-\frac{\pi}{8}(r\Delta t)^{2}\right)}}$
      $\bar{\gamma}_{*}$ $:=\frac{2\left(1-\exp\left(-\frac{1}{2}(r+\frac{\bar{\sigma}_{*}^{2}}{4})\Delta t\right)\right)}{1-\exp\left(-\frac{1}{2}(r+\frac{\bar{\sigma}_{*}^{2}}{4})\Delta t\right)+\Phi\left(\frac{(r+\frac{\bar{\sigma}_{*}^{2}}{2})\sqrt{\Delta t}}{\bar{\sigma}_{*}}\right)-e^{-r\Delta t}\Phi\left(\frac{(r-\frac{\bar{\sigma}_{*}^{2}}{2})\sqrt{\Delta t}}{\bar{\sigma}_{*}}\right)}.$
        If $\hat{\gamma}<\hat{\gamma}^{*}(\bar{\sigma})$ then there does not exist a calibrated volatility.
        If $\hat{\gamma}=\hat{\gamma}^{*}(\bar{\sigma})$ then:
          there does not exist a calibrated volatility if $G_{C}(\bar{\sigma})\neq 0$; and
          there exists a unique calibrated volatility $\sigma^{M}=\bar{\sigma}$ if $G_{C}(\bar{\sigma})=0$.
        If $\hat{\gamma}>\hat{\gamma}^{*}(\bar{\sigma})$ and $\gamma\leq\bar{\gamma}_{*}$ then:
          there does not exist a calibrated volatility if $G_{C}(\sigma^{I}_{1})G_{C}(\sigma^{I}_{2})>0$; and
          there exists a unique calibrated volatility $\sigma^{M}\in[\sigma^{I}_{1},\sigma^{I}_{2}]$ if $G_{C}(\sigma^{I}_{1})G_{C}(\sigma^{I}_{2})\leq 0$.
        If $\hat{\gamma}>\hat{\gamma}^{*}(\bar{\sigma})$ and $\gamma>\bar{\gamma}_{*}$ then:
          there does not exist a calibrated volatility if $G_{C}(\bar{\sigma}_{*})>0$;
          there exists a unique calibrated volatility $\sigma^{M}\in[\sigma^{I}_{1},\bar{\sigma}_{*}]$ if $G_{C}(\sigma^{I}_{1})G_{C}(\bar{\sigma}_{*})\leq 0$; and
          there exists a unique calibrated volatility $\sigma^{M}\in[\bar{\sigma}_{*},\sigma^{I}_{2}]$ if $G_{C}(\sigma^{I}_{2})G_{C}(\bar{\sigma}_{*})\leq 0$.
        We wish to note that these final two cases can both simultaneously be satisfied leading to the possibility of two distinct calibrated volatiles (similarly, it is possible that neither is satisfied with $G_{C}(\bar{\sigma}_{*})<0$ so that no calibrated volatility exists).
Proof.
We wish to note that (4.1) is satisfied at $\sigma^{M}>0$ if and only if $G_{C}(\sigma^{M})=0$. As it will be used throughout this proof, we provide here the derivative $G_{C}^{\prime}(\sigma)=\operatorname{e}^{-\frac{1}{2}\left(r+\frac{\sigma^{2}}{4}\right)\Delta t}\left[\frac{\sigma\Delta t}{4}-\sqrt{\frac{\Delta t}{2\pi}}e^{-\frac{r^{2}\Delta t}{2\sigma^{2}}}\right]$ for any $\sigma>0$. Finally, for any implied volatility $\sigma^{I}>0$, $G_{C}(\sigma^{I})=C-\frac{2}{\hat{\gamma}}\left(1-\operatorname{e}^{-\frac{1}{2}(r+\frac{(\sigma^{I})^{2}}{2})\Delta t}\right)$.
  • Let $r=0$. Trivially $G_{C}^{\prime}(\bar{\sigma}_{*}^{0})=0$ with $G_{C}^{\prime}(\sigma)<0$ for any $\sigma<\bar{\sigma}_{*}^{0}$ and $G_{C}^{\prime}(\sigma)>0$ for any $\sigma>\bar{\sigma}_{*}^{0}$.
    • Assuming $\gamma\leq\bar{\gamma}_{*}^{0}$, the unique implied volatility satisfies $\sigma^{I}\leq\bar{\sigma}_{*}$. Therefore, noting that $\lim_{\sigma\searrow 0}G_{C}(\sigma)=C>0$, there exists at most one calibrated volatility in $(0,\sigma^{I}]$ with existence if and only if $G_{C}(\sigma^{I})\leq 0$.
    • Assuming $\gamma>\bar{\gamma}_{*}^{0}$, the unique implied volatility satisfies $\sigma^{I}>\bar{\sigma}_{*}$. Therefore, there are between 0 and 2 possible calibrated volatiles. Using $\lim_{\sigma\searrow 0}G_{C}(\sigma)=C>0$ and the construction of $\bar{\sigma}_{*}^{0}$, the stated cases follow.
  • Let $r>0$. Trivially $G_{C}^{\prime}(\bar{\sigma}_{*})=0$ with $G_{C}^{\prime}(\sigma)<0$ for any $\sigma<\bar{\sigma}_{*}$ and $G_{C}^{\prime}(\sigma)>0$ for any $\sigma>\bar{\sigma}_{*}$. We will consider only the cases with $\hat{\gamma}>\hat{\gamma}^{*}(\bar{\sigma})$ as all others trivially hold.
    • Assuming $\gamma\leq\bar{\gamma}_{*}$, then it can be verified that $\bar{\sigma}_{*}\not\in(\sigma^{I}_{1},\sigma^{I}_{2})$. Therefore, $G_{C}(\sigma)$ is monotonic on $\sigma\in[\sigma^{I}_{1},\sigma^{I}_{2}]$. As a direct consequence, the stated cases follow as provided.
    • Assuming $\gamma>\bar{\gamma}_{*}$, then it can be verified that $\bar{\sigma}_{*}\in(\sigma^{I}_{1},\sigma^{I}_{2})$. As, by construction, $G_{C}(\sigma)$ reaches its minimum at $\bar{\sigma}_{*}$, there exists a root to $G_{C}$ if and only if $G_{C}(\bar{\sigma}_{*})\leq 0$. Using monotonicity of $G_{C}$ on $[\sigma^{I}_{1},\bar{\sigma}_{*}]$ and $[\bar{\sigma}_{*},\sigma^{I}_{2}]$, the conditions follow.
We conclude this section by revisiting Example 2.3 to calibrate the volatility to empirical data. Within this example we find that the proposed procedure can accurately re-price the liquidity token so as to effectively eliminate the arbitrage opportunities encountered in practice based on the prevailing market price of $2\sqrt{P}$.
@@
Consider the USDC/WETH Uniswap v3 pool on the Polygon blockchain considered in Example 2.3 (i.e., $\Delta t=2$ seconds and $\gamma=5bps$, i.e., $\hat{\gamma}\approx 5.0025bps$). As in Example 2.3, throughout this discussion we will set $r=5\%$. Recall, also, from the third case in Example 4.4 there exist two distinct implied volatilities for the market price of $2\sqrt{P}$ given by $\sigma^{I}_{1}\approx 6.44\%$ and $\sigma^{I}_{2}\approx 310.47\%$ (annualized) since $\Delta t<\overline{\Delta t}\approx 42.40$ hours. Using the USDC/WETH data of Example 2.3, we can calibrate $C\approx 2.5937\times 10^{-5}$. [^8] So as to more accurately approximate the geometric Brownian motion, the calibrated $C$ only considers blocks in which a swap occurred, i.e., so that $N$ is the number of blocks with swaps rather than the total number of blocks within June 2023. To determine which case of Corollary 4.5 we fall into, we note that $\gamma<\bar{\gamma}_{*}\approx 64.32\%$ and $\bar{\sigma}_{*}\approx 6,336.63\%>\sigma^{I}_{2}$. Noting that $G_{C}(\sigma^{I}_{1})\approx 1.95\times 10^{-5}>0>G_{C}(\sigma^{I}_{2})\approx-2.86\times 10^{-4}$, there exists a unique volatility $\sigma^{M}\approx 25.82\%$ (annualized) which fits the observed Uniswap data which admits an arbitrage at the market price of $2\sqrt{P}$ (as observed in Example 2.3). Finally, with this approximation of the calibrated volatility, we are able to re-price the liquidity token by noting that $\hat{\gamma}/\hat{\gamma}^{*}(\sigma^{M})\approx 3.069>1$, i.e., Uniswap is underpricing the liquidity token by a factor of $3.069$.
In Figure 5, we compare the delta hedged position under this re-pricing (solid blue line) compared to that of the original market price (dashed black line); for direct comparisons, we assume 1 USDC was invested in the pool for both cases (yielding different number of liquidity tokens). Notably, the hedging error under repricing is an order of magnitude lower and fluctuates around the initial price of 1 compared to the market pricing (as observed already in Example 2.3). This minimal hedging error demonstrates that this updated pricing can be viewed as an arbitrage-free price of the liquidity token. [^9] Despite the nearly perfect hedge constructed for the re-priced CPMM, we note that there exist a select few times when its value jumps. These jumps correspond to times at which the WETH price experiences large movements, see Figure 1(a). Also, recall that we are using a constant calibrated volatility $\sigma^{M}$ throughout the period of study rather than allowing it to fluctuate; alternative AMM constructions which permit variable implied volatilities are discussed in Section 5.
Figure 5: Example 4.6: Comparison of the delta hedged position of 1 USDC investment in a liquidity token under market pricing (black dashed line) and with the calibrated arbitrage-free price (blue solid line).
@@
Herein we have focused on hedging the liquidity token itself rather than the impermanent loss as in, e.g., [11, 17]. As the impermanent loss is just the difference between the value of the liquidity token (including any collected fees) and the value of the initial position $P_{t}x_{0}+e^{rt}y_{0}=P_{t}/\sqrt{P_{0}}+e^{rt}\sqrt{P_{0}}$ which was used to mint the liquidity token. [^10] We wish to note that this construction may need adjustments if the price of the liquidity token grows beyond $2\sqrt{P_{0}}$ at time 00 as we propose in Example 4.6 above. As the initial position is static, hedging the impermanent loss reduces to appropriately hedging the liquidity token itself and, therefore, our results easily generalize to this more widely studied problem.

### v3

@@ 4.1 Implied Volatility
First, consider the case with zero risk-free rate $r=0$. We note that $\lim_{\sigma\searrow 0}G(\sigma)=0$ (with $\lim_{\sigma\searrow 0}G^{\prime}(\sigma)=-\hat{\gamma}\sqrt{\frac{\Delta t}{2\pi}}$) and $\lim_{\sigma\nearrow\infty}G(\sigma)=2$. Therefore, there exists at least one implied volatility $\sigma^{I}>0$. Additionally, $\bar{\sigma}:=\frac{\hat{\gamma}}{2+\hat{\gamma}}\sqrt{\frac{8}{\pi\Delta t}}$ is the unique positive root of $G^{\prime}$. Therefore, there cannot exist a positive implied volatility below $\bar{\sigma}$ and there must exist a unique implied volatility $\sigma^{I}>\bar{\sigma}$.
Assume, now, a strictly positive risk-free rate $r>0$. First, we note that $\lim_{\sigma\searrow 0}G(\sigma)=2(1-e^{-\frac{1}{2}r\Delta t})-\hat{\gamma}(e^{-\frac{1}{2}r\Delta t}-e^{-r\Delta t})>0$ and $\lim_{\sigma\nearrow\infty}G(\sigma)=2$. Therefore, there exists an implied volatility $\sigma^{I}>0$ if, and only if, $\inf_{\sigma>0}G(\sigma)\leq 0$. To investigate this infimum, we first consider the sign of $G^{\prime}$. In particular, $G^{\prime}(\sigma)\leq 0$ if, and only if, $(2+\hat{\gamma})\frac{\sigma\Delta t}{4}-\hat{\gamma}\sqrt{\frac{\Delta t}{2\pi}}e^{-\frac{r^{2}\Delta t}{2\sigma^{2}}}\leq 0$ or, equivalently via rearranging terms, $\sigma e^{\frac{r^{2}\Delta t}{2\sigma^{2}}}\leq\frac{4\hat{\gamma}}{\sqrt{2\pi\Delta t}(2+\hat{\gamma})}$. Noting that $\inf_{\sigma>0}\sigma e^{\frac{r^{2}\Delta t}{2\sigma^{2}}}=r\sqrt{\Delta t}e^{1/2}$, $G^{\prime}(\sigma)>0$ for every $\sigma>0$ if, and only if, $r\sqrt{\Delta t}e^{1/2}>\frac{4\hat{\gamma}}{\sqrt{2\pi\Delta t}(2+\hat{\gamma})}$, i.e., $\Delta t>\overline{\Delta t}$. Assuming $\Delta t\leq\overline{\Delta t}$, we can find a unique root of $G^{\prime}$ using the Lambert $W$ function at $\bar{\sigma}$ as provided in the statement of the lemma. We note that $\Delta t\leq\overline{\Delta t}$ guarantees $\frac{\pi}{2}\left[\frac{(2+\hat{\gamma})r\Delta t}{2\hat{\gamma}}\right]^{2}\in[-e^{-1},0)$, i.e., $\bar{\sigma}$ is well-defined and positive. Assume $\Delta t\leq\overline{\Delta t}$, then since this root $\bar{\sigma}$ is unique, it must follow that $\inf_{\sigma>0}G(\sigma)\in\{G(\bar{\sigma}),\lim_{\sigma\searrow 0}G(\sigma)\}$. From this, and recalling the relation between the sign of $G$ and the risk-neutral valuation, the result trivially follows. ∎
@@
Before continuing, we want to provide a numerical example to demonstrate all possible outcomes for the set of implied volatilities. We will do this by considering the Polygon blockchain with different fee levels $\gamma$ to demonstrate the different possible settings for the implied volatility. In each of these cases, $\Delta t<\overline{\Delta t}$ by orders of magnitude.
@@
Consider a constant product market maker on the Polygon blockchain ($\Delta t=2$ seconds). For these examples, recall that the risk-neutral valuation and other formulas employed throughout this work utilize the realized fee level $\hat{\gamma}=\gamma/(1-\gamma)$. For any choice of $\gamma$, there are two possible outcomes. The less interesting possibility is that $V_{0}(P)<2\sqrt{P}$ so that depositing liquidity for the price of $2\sqrt{P}$ is expected to lose value, i.e., the expected discounted value of the fees would not cover this initial deposit value. [^4] Within this work, we have assumed that it is not possible to sell liquidity tokens short and, therefore, no arbitrage opportunity exists if $V_{0}(P)<2\sqrt{P}$. The more interesting scenario is where $V_{0}(P)\geq 2\sqrt{P}$ which (with a strict inequality) results in an arbitrage opportunity since the purchase price of the liquidity tokens $2\sqrt{P}$ is below the risk-neutral valuation of the fees. Therefore, in this latter scenario, a rational investor can deposit the liquidity for $2\sqrt{P}$ and, with appropriate hedging, obtain the value $V_{0}(P)$. [^5] We refer the interested reader to [16, Chapter 1.4] for an initial discussion on arbitrage.
(a) $r=0\%$ with marked point indicating the implied volatility $\sigma^{I}$ at $\gamma=5bps$.
(b) $r=5\%$ (annualized) with marked points indicating the implied volatilities $\sigma^{I}$ at $\gamma\in\{1bps,1.14114bps,5bps\}$.
@@
Figure 3: Example 4.4: The shaded region indicates fee-volatility $(\gamma,\sigma)$ pairs that provide arbitrage opportunities.
@@
  • If $r=0\%$ and $\gamma=5bps$ then, by Lemma 4.2, there exists a unique implied volatility $\sigma^{I}\approx 316.75\%$ (annualized). Notably, as proven in Corollary 4.3 and shown in Figure 3(a), if $\sigma<\sigma^{I}$ then $V(P_{0})>2\sqrt{P_{0}}$ and a risk-neutral investor would deposit liquidity at the AMM. In contrast, if $\sigma>\sigma^{I}$, then $V(P_{0})<2\sqrt{P_{0}}$ and a risk-neutral investor would not pool liquidity at the AMM.
  • If $r=5\%$ (annualized) and $\gamma=1bps$ then $\overline{\Delta t}\approx 8.48$ hours and $\bar{\sigma}\approx 0.3168$ with $\hat{\gamma}^{*}(\bar{\sigma})\approx 1.4962bps>\hat{\gamma}$. Therefore, there does not exist an implied volatility at this fee level, i.e., a risk-neutral investor would never deposit liquidity at the AMM. This is shown in Figure 3(b) by the dotted line.
  • If $r=5\%$ (annualized) and $\gamma\approx 1.4114bps$ (so that $\hat{\gamma}\approx 1.4116bps$) then $\overline{\Delta t}\approx 11.97$ hours and $\bar{\sigma}\approx 0.4472$ with $\hat{\gamma}^{*}(\bar{\sigma})=\hat{\gamma}$. Therefore, the unique implied volatility is given by $\sigma^{I}=\bar{\sigma}\approx 44.72\%$ (annualized). Notably, if $\sigma\neq\sigma^{I}$ then a risk-neutral investor would not pool liquidity at the AMM.
  • If $r=5\%$ (annualized) and $\gamma=5bps$ then $\overline{\Delta t}\approx 42.40$ hours and $\bar{\sigma}\approx 1.5846$ with $\hat{\gamma}^{*}(\bar{\sigma})\approx 2.7002bps<\hat{\gamma}$. Therefore, there exists two implied volatilities: $\sigma^{I}_{1}\approx 6.44<\bar{\sigma}<\sigma^{I}_{2}\approx 310.47\%$ (annualized). Notably, as proven in Corollary 4.3 and shown Figure 3(b), if $\sigma\in(\sigma^{I}_{1},\sigma^{I}_{2})$ then $V(P_{0})>2\sqrt{P_{0}}$ and a risk-neutral investor would deposit liquidity at the AMM. In contrast, if $\sigma\in(0,\sigma^{I}_{1})\cup(\sigma^{I}_{2},\infty)$ then $V(P_{0})<2\sqrt{P_{0}}$ and a risk-neutral investor would not pool liquidity at the AMM.
@@
As evidenced in Example 4.4 above (and comparing to the parameters of, e.g., Example 2.3), non-uniqueness of the implied volatility can easily occur in practice when $r>0$. It becomes important to understand which of $\{\sigma^{I}_{1},\sigma^{I}_{2}\}$ should be quoted. Herein, as $\sigma^{I}_{2}$ converges to the unique solution $\sigma^{I}$ when the risk-free rate approaches 0, i.e., $\lim_{r\searrow 0}\sigma^{I}_{2}(r)=\sigma^{I}(0)$ with dependence on the risk-free rate made explicit, we take this upper implied volatility to be the more meaningful setting. In comparison, the lower implied volatility $\sigma^{I}_{1}$ converges to 0 as the risk-free rate approaches 0, i.e., $\lim_{r\searrow 0}\sigma^{I}_{1}(r)=0$.
@@ 4.2 Estimating the Arbitrage-Free Price of a Liquidity Token
In contrast to the market price $2\sqrt{P}$ utilized for the implied volatility in Section 4.1 above, herein we want to calibrate the pricing of the liquidity token to the observed data. To do this, we will first find the volatility $\sigma^{M}>0$ so that we observe martingale pricing in the data. Specifically, over a given historical time period (e.g., 1 month), we find the volatility $\sigma^{M}>0$ so that the delta hedged position under pricing with $\hat{\gamma}^{*}(\sigma^{M})$ at the end of the period has the same value as the start of the period. We note that this historically calibrated volatility differs from the more traditional historical volatility based on repeated snapshots of the data. To demonstrate the long-term stability of this method, we revisit Example 2.3 to calibrate the volatility to empirical data. Within this example, we find that the proposed method can accurately re-price the liquidity token so as to effectively eliminate the arbitrage opportunities encountered in practice based on the prevailing market price of $2\sqrt{P}$ even months after the calibration period.
@@
Consider the USDC/WETH Uniswap v3 pool on the Polygon blockchain considered in Example 2.3 (i.e., $\Delta t=2$ seconds and $\gamma=5bps$, i.e., $\hat{\gamma}\approx 5.0025bps$). As in Example 2.3, throughout this discussion we will set $r=0\%$. Recall, also, from the first case in Example 4.4 there exist a unique implied volatilities for the market price of $2\sqrt{P}$ given by $\sigma^{I}\approx 316.75\%$ (annualized). Specifically, calibrating the volatility $\sigma^{M}>0$ to January 2023, we find $\sigma^{M}\approx 143.75\%$ (annualized) which we then apply throughout the full 2023 calendar year. We note that, with this approximation of the calibrated volatility, we are able to re-price the liquidity token by noting that $\hat{\gamma}/\hat{\gamma}^{*}(\sigma^{M})\approx 2.2048>1$, i.e., Uniswap is underpricing the liquidity token by a factor of $2.2048$.
In Figure 4, we compare the delta hedged position under this re-pricing (solid black line) compared to that of the original market price (dashed blue line); for direct comparisons, we assume 1 USDC was invested in the pool for both cases (yielding different number of liquidity tokens). Notably, the hedging error under repricing is an order of magnitude lower and fluctuates around the initial price of 1 compared to the market pricing (as observed already in Example 2.3). This minimal hedging error demonstrates that this updated pricing can be viewed as an arbitrage-free price of the liquidity token. Despite the nearly perfect hedge constructed for the re-priced CPMM, we note that there exist a select few times when its value jumps. These jumps correspond to times at which the WETH price experiences sudden, large movements which are not delta hedgeable, see Figure 1(a). As exected, the power of this repricing decreases as we move further into our out-of-sample period at the end of 2023 with a noticeable updward trend from October through December; recall we are using a constant calibrated volatility $\sigma^{M}$ based on only January’s data rather than allowing it to fluctuate throughout the period. Alternative AMM constructions which permit variable implied volatilities are discussed in Section 5.
Figure 4: Example 4.5: Comparison of the delta hedged position of 1 USDC investment in a liquidity token under market pricing (blue dashed line) and with the calibrated arbitrage-free price (black solid line). Volatility is calibrated on January 2023 and applied throughout calendar year 2023.
@@
Herein we have focused on hedging the liquidity token itself rather than the impermanent loss as in, e.g., [13, 18]. As the impermanent loss is just the difference between the value of the liquidity token (including any collected fees) and the value of the initial position $P_{t}x_{0}+e^{rt}y_{0}=P_{t}/\sqrt{P_{0}}+e^{rt}\sqrt{P_{0}}$ which was used to mint the liquidity token. [^6] We wish to note that this construction may need adjustments if the price of the liquidity token grows beyond $2\sqrt{P_{0}}$ at time 0 as we propose in Example 4.5 above. As the initial position is static, hedging the impermanent loss reduces to appropriately hedging the liquidity token itself and, therefore, our results easily generalize to this more widely studied problem.